# Omni-View: Unlocking How Generation Facilitates Understanding in Unified 3D Model based on Multiview images

### v1

@@ 4.3 Ablation studies
Effect of the D2S (dense-to-sparse). The efficacy of D2S training is substantiated during stage 1. The results show that D2S significantly improves the model understanding performance, exceeding the visual reconstruction method advanced in Ross3D (wang2025ross3d) for scene understanding tasks necessitating spatiotemporal ordering. In this ablation, the geometry module is kept trainable.
@@
Table 6: Ablation on the D2S mechanism in stage 1 (S1). “None” means we don’t train generation model in this experiment. The “random mask” denotes the visual reconstruction method in Ross3D (wang2025ross3d).
@@ 5 Conclusion
We introduce Omni-View, a unified 3D scene understanding and generation model. By decomposing the generation model into distinct texture and geometry components, we establish the feasibility and efficacy of employing generation processes to enhance 3D scene understanding and spatial reasoning. Our method illustrates that a unified understanding and generation model is capable of achieving performance on par with leading specialized understanding models. We posit that Omni-View can serve as a foundational model across 3D and multiview domains, thereby advancing the development of downstream applications such as spatial and intelligence.
@@
Reproducibility Statement. We state that Omni-View is highly reproducible. Implementation details on our main experiences are provided in Section 4.1. It is anticipated that these descriptions can sufficiently demonstrate the reproducibility of Omni-View. We plan to open-source the code and weight files after the paper passes peer review.
Limitations and future work. Although Omni-View has demonstrated a unified advance in the domains of question answering, spatial reasoning, and novel view synthesis, its grounding capabilities remain to be substantiated. Additionally, its generation model currently lacks the capability for long-range world generation. Future efforts will be concentrated on using reinforcement learning to augment Omni-View’s performance in 3D visual grounding and long-range generation. At the same time, since the training data used by its geometry module is a synthetic depth map, its actual geometry prediction ability may not be accurate enough.
Appendix A Qualitative results.
@@ A.2 Spatial Reasoning
A.3 NVS from single view
We show an effective result (3f2bf7371b72e40f) and a bad case (e7020d449f50c737) for NVS from single view task. It can be seen that when the angle changes greatly, Omni-View’s inter-frame consistency still has room for improvement, for example, the object on the left side of the door is inconsistent between View-20 and View-24 in sample e7020d449f50c737.
Reference (3f2bf7371b72e40f)
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@ A.3 NVS from single view
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Reference (e7020d449f50c737)
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
The Use of Large Language Models (LLMs). LLMs are used to correct potential grammatical inaccuracies in the manuscript. LLMs do not participate in research ideation.

### v2

@@ 4.3 Ablation studies
Effect of the D2S (dense-to-sparse). The efficacy of D2S training is substantiated during stage 1. The results show that D2S significantly improves the model understanding performance, exceeding the visual reconstruction method advanced in Ross3D (Wang et al., 2025a) for scene understanding tasks necessitating spatiotemporal ordering. In this ablation, the geometry module is kept trainable.
@@
Table 6: Ablation on the D2S mechanism in stage 1 (S1). “None” means we don’t train generation model in this experiment. The “random mask” denotes the visual reconstruction method in Ross3D (Wang et al., 2025a).
@@ 5 Conclusion
We introduce Omni-View, a unified 3D scene understanding and generation model. By decomposing the generation model into distinct texture and geometry components, we establish the feasibility and efficacy of employing generation processes to enhance 3D scene understanding and spatial reasoning. Our method illustrates that a unified understanding and generation model is capable of achieving performance on par with the leading specialized understanding models. We posit that Omni-View can serve as a foundational model across 3D and multiview domains, thereby advancing the development of downstream applications such as spatial intelligence.
@@
Reproducibility Statement. We state that Omni-View is highly reproducible. Implementation details on our main experiences are provided in Section 4.1 and Appendix B. It is anticipated that these descriptions can sufficiently demonstrate the reproducibility of Omni-View. We plan to open-source the code and weight files after the paper passes peer review.
Appendix A Qualitative results
@@ A.2 Spatial Reasoning
A.3 Scale consistency across scenes
We analyze the scale consistency across different scenes. To this end, we selected some test samples from the SPAR-7M dataset and used Omni-View to perform absolute distance prediction based on multi-view images. This SPAR-7M consists of scenes from three datasets: ScanNet and ScanNet++, where the average scene depth ranges from 3 to 5 meters with a maximum depth of 12 meters; and Structured3D, which features larger-scale environments with average depths between 4 and 6 meters and maximum depths reaching up to 20 meters. Qualitative visualizations demonstrate that Omni-View can predict absolute depth in scenes with different metric scale.
A.4 NVS from single view
Indoor scenes. We present the two visual results of NVS in indoor scenes. Omni-View can reasonably imagine unseen areas while maintaining the texture and structure of observed objects, like the followers on the table in the first row.
Reference View-12 View-24 View-36 View-48
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Outdoor scenes. We present the two visual results of NVS in outdoor scenes. When the camera movement is small, Omni-View can consistently generate new views.
Reference View-8 View-16 View-24 View-32
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Depth estimation. We present the depth prediction results of Omni-View in indoor and outdoor scenes. In indoor scenes, the output of Omni-View’s geometry module is more accurate.
Reference View-10 View-20 View-30 View-40
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Failure case. We claim that the existing Omni-View is inadequate for effectively processing outdoor scenes with substantial camera movement. As illustrated in the figure below, the result images from Omni-View exhibit significant artifacts, highlighted within the red dashed boxes. Future work will focus on resolving the following challenges to enhance the handling of outdoor scenes with extensive camera motion: (1) the development of more precise camera control mechanisms, and (2) the improvement of inter-frame texture consistency stability.
Reference View-2 View-4 View-6 View-8
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
A.5 Effect of generation in grid
Generation in grid. We explored small improvements to enhance inter-frame consistency and support long-sequence scene generation: “generation in grid”. Specifically, we arrange 4 consecutive views into a single frame by organizing them in a grid layout, and perform autoregressive generation over the resulting sequence of grid-organized frames.
Figure 2: Generation in grid.
Results. As shown in the figures on the next page, we demonstrate that after using “generation in grid”, Omni-View can generate more consistent novel view images.
Reference “Generation in grid”
Refer to caption Refer to caption Refer to caption
@@ A.5 Effect of generation in grid
Refer to caption Refer to caption Refer to caption
@@
Reference “Generation in frame-by-frame”
Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption
@@
Figure 3: Different inference strategy in NVS (RealEstate10k, e7020d449f50c737). The first three rows use the generation in grid strategy, achieving significant improvements in inter-frame consistency. The last three rows still use the frame-by-frame generation strategy, which has poor inter-frame consistency.
Appendix B Technical details
B.1 Dataset
3D scene understanding and spatial reasoning. We curate a filtered training dataset containing 780k valid samples by consolidating data from multiple sources, including SQA3D, ScanQA, 3DOD, ScanRefer, VLM-3R, a 234k subset of SPAR from VG-LLM, and a 64k subset of llava-hound4 from VG-LLM. To ensure data quality, we perform two main filtering steps: (i) deduplication across the combined dataset, and (ii) removal of samples with invalid bounding box annotations. Specifically, we convert all bounding boxes to the $x_{1},y_{1},x_{2},y_{2}]$ format and exclude any instance where $x_{1}<0$, $y_{1}<0$, $x_{2}>\text{width}$, or $y_{2}>\text{height}$.
Novel view synthesis. We select 61k video clips from RealE10K train set, excluding those with fewer than 32 frames or significant motion blur that could degrade training stability. Depth maps are synthesized using the Voyager data pipeline, while captions are generated via QwenVLMax.
B.2 Architecture.
Understanding model. The detailed architecture is shown in Figure 4. The text tokenizer inherits from the vocabulary used by Qwen2. The image tokenizer for understanding is SigLIP. The backbone has layers as 28, attention head as 28, and hidden size as 3584, totaling 7B parameters. Key architectural choices include SwiGLU as the activation function, RMSNorm for normalization, FlexAttention (via PyTorch) for efficient self-attention computation, and position encoding adapted from Qwen2.
Figure 4: Architecture of understanding model and texture module.
Texture module. This module uses FLUX-VAE as the image tokenizer, sharing the same backbone architecture as the understanding model.
Figure 5: Architecture of geometry module in two stages.
Geometry module. The detailed architecture is shown in Figure 5. Depth maps are also encoded by FLUX-VAE. Its backbone is basically the same as the block architecture, but reduced to 4 layers, containing about 1B parameters. Unlike the others, the architecture of the geometry module changes across different training stages. In stage 1, it receives features from the understanding model for cross-attention. In stage 2, it no longer receives outputs from the understanding model. The reason for this design can be found in the discussion of stage 2 in Section 3.2.
B.3 Training and computational cost.
All training phases use the AdamW optimizer with $\beta_{1}=0.9,\beta_{2}=0.95$, a peak learning rate of $1\times 10^{-5}$, and a linear warm-up schedule covering the first 5% of iterations.
Stage 1. We train on the 3D scene understanding dataset for 10,000 iterations with a packed sequence length of 50k, using 32 H100 GPUs, which takes approximately 160 hours.
Stage 2. For the generation task, we perform 20,000 iterations on the novel view synthesis dataset with a packed sequence length of 32k, using 32 H100 GPUs, requiring approximately 40 hours in total.
Understanding inference. The model takes approximately 2.5 seconds on average to process a 32-frame multi-view scene understanding query using a single H100 GPU.
Generation inference. Generating a single image at resolution 640 $\times$ 352 takes about 2.2 seconds on average with one H100 GPU.
B.4 Convergence behavior.
Stage 1. Overall, most losses exhibit smooth and stable convergence. However, we observe spikes in the camera pose loss during training, particularly in later epochs. We hypothesize that this behavior may stem from instable optimizion by learnable queries in the camera pose estimation task. However, these fluctuations do not prevent the model from converging to effective solutions, as evidenced by strong performance in 3D scene understanding, spatial reasoning, and novel view generation.
Stage 2. Since we used a grid generation format in stage 2, the per-image prediction method in stage 1 cannot be directly transferred to the “generation in grid” paradigm in stage 2. However, due to the generative prior learned in stage 1, Omni-View can quickly converge to a lower loss in stage 2. However, it was observed that the geometry loss showed a slight increase at around 8000 iterations, followed by a decrease, suggesting that it might be possible to further reduce $\lambda_{geo}$ to improve the training stability.
B.5 Activation visualization
We visualize the activation maps of Bagel and Omni-View when they perform spatial reasoning tasks, as shown in the figure below. In this example, we want the model to locate how many cabinets are in the current scene, using the prompt: “How many cabinet(s) are in this room?”. It can be seen that Bagel mainly focuses on the first two images that are irrelevant to the question, whereas Omni-View is able to attend to each image as much as possible. The wider attention may be the reason why Omni-View performs better on 3D scene understanding tasks.
Bagel-FT
Omni-View (Ours)
Figure 6: Activation map. Bagel-FT vs. Omni-View.
Appendix C The Use of Large Language Models (LLMs)
LLMs are used to correct potential grammatical inaccuracies in the manuscript. LLMs do not participate in research ideation.